# Optimal and Bounded Suboptimal Any-Angle Multi-agent Pathfinding

### v1

@@ VI Experimental Results
We implemented AA-CCBS in C++ [^1] https://github.com/PathPlanning/Continuous-CBS/tree/AA-CCBS and conducted the experimental evaluation on five different maps from the well-known in the community MovingAI benchmark [8]: empty-16-16, random-32-32-20, maze-32-32-4, den321d and warehouse-10-20-10-2-2. For each map, the benchmark provides 25 scenarios. Each scenario is a set of start-goal locations. We take two agents, run the algorithm, and then increment the number of agents until the algorithm is not able to find a solution within a time-limit of 300 seconds. In the latter case, we proceed to the next scenario. We evaluate six versions of AA-CCBS: AA-CCBS (vanilla), AA-CCBS+MC1, AA-CCBS+MC2, AA-CCBS+MC3, AA-CCBS+DS, AA-CCBS+DS+MC3.
Fig. 4 shows the results. The left pane shows the number of solved instances (with the breakdowns for each map). The results indicate that MC1 and MC2 do not provide significant improvement over vanilla AA-CCBS. MC3, DS, and DS+MC3, however, notably outperform it, especially the ones that use DS. The central pane shows the number of tasks solved within a certain time cap. The X-axis shows the number of tasks and Y-axis shows the time needed to solve this amount of tasks. One can note that MC3 outperforms all other variants of AA-CCBS for easy tasks, i.e. for the ones that need less than 0.1 seconds to be solved. Meanwhile DS variants of AA-CCBS clearly outperform MC3 when the time, needed to solve an instance, is greater than 1 second. The right pane of Fig. 4 depicts the number of the high-level iterations of AA-CCBS with MC3, DS, DS+MC3 for each MAPF instance that was solved by all these methods. As seen on the plot, this number is lower for MC3 for nearly half of the instances. This also confirms that MC3 solves some (i.e. easy) instances faster than DS and MC2+DS, while it is outperformed on the other (hard) ones.
Overall the obtained results confirm that both MC and DS are valuable techniques to enhance AA-CCBS performance. MCs (specifically, MC3) are well-suited for the easy instances while the DS – for the harder ones. Generally, we would recommend to use AA-CCBS+DS+MC3 to solve any-angle MAPF optimally.
@@ VII Bounded-Suboptimal AA-CCBS
Given a real number $w>1$ a bounded suboptimal (BS) solution is the one whose cost does not exceed the cost of the optimal one by more than a factor of $w$. Finding a BS solution is often much easier and faster than finding an optimal one. Creating a BS variant of AA-CCBS with any of our enhancements is straightforward: replacing the best-first search in the AA-CCBS high-level with a Focal search [33]. [^2] Creating a BS low-level search is not trivial in our case and goes beyond the scope of this paper. I.e. at each high-level iteration we, first, form a subset of nodes whose cost does not exceed the cost of the most promising node (the one which regular AA-CCBS would have picked) by a factor of $w$ and, second, choose from this subset the node with the smallest number of conflicts breaking ties in favor of nodes containing more constraints. Thus we force the search to more greedily resolve the conflicts, while still maintaining the bound on the suboptimality of the resultant solution.
We implemented such a BS version of AA-CCBS with the three best-performing enhancements, DS, MC3, DS+MC3, and evaluated them on the same problem instances as before. Fig. 5 shows the results. In the upper part the box-and-whiskers plot of the resultant solution cost depending on the suboptimality bound are shown. The table at the bottom shows the number of the solved instances across all maps.
Indeed, MC3 notably increases the performance of AA-CCBS when searching for bounded suboptimal solutions (i.e. when $w>1$), while DS variants are not even able to outperform vanilla AA-CCBS. We hypothesize that this is due to the positive constraints imposed by DS do not lead to immediate eliminating the conflict that causes these constraints. When searching for optimal solutions this is compensated by reducing the branching factor. However the search for suboptimal solutions is more greedy by design and branching factor becomes much less important. Thus, DS looses much of its power, while retaining the drawback of not immediately eliminating the conflict. At the same time, MC3 imposes constraints on a larger number of actions, and as the search greedily goes deeper it actually eliminates the conflict. Moreover, due to the requirement to satisfy all the imposed positive constraints, the actual costs of the solutions found by DS versions are on average more than twice higher than the ones of MC3. The actual overhead in solution costs of MC3 version is not higher than 2% on average even when the suboptimality bound is $w=1.25$. The obtained results provide further support that MC3 is a useful technique: not only does it speed the search for optimal solutions on easy instances, but it also significantly increases the performance of bounded-suboptimal AA-CCBS.

### v2

@@ VI Experimental Results
We implemented AA-CCBS in C++ [^1] https://github.com/PathPlanning/Continuous-CBS/tree/AA-CCBS and conducted the evaluation on different maps from the MovingAI benchmark [8]: empty-16-16, random-32-32-20, maze-32-32-4, den321d and warehouse-10-20-10-2-2. For each map, the benchmark provides 25 scenarios. Each scenario is a set of start-goal locations. We take two agents, run the algorithm, and then increment the number of agents until the algorithm is not able to find a solution within a time limit of 300 seconds. In the latter case, we proceed to the next scenario. We evaluate six versions of AA-CCBS: AA-CCBS (vanilla), AA-CCBS+MC1, AA-CCBS+MC2, AA-CCBS+MC3, AA-CCBS+DS, AA-CCBS+DS+MC3.
Fig. 4 shows the results. The left pane shows the coverage, which is the number of solved instances within our 300 seconds time limit. The colors show the breakdowns for each map. The results indicate that MC1 and MC2 do not provide a significant improvement over vanilla AA-CCBS, while MC3, and especially DS and DS+MC3, outperform it significantly. The central pane provides a finer-grained analysis, plotting the number of instances solved ($x$-axis) for different time limits ($y$-axis). We filtered out trivial instances where the initial any-angle paths of the agents do not contain any conflict. This analysis shows that MC3 outperforms all other variants of AA-CCBS if the time limit is between 0.01 and 0.1 seconds, while DS and DS+MC3 outperform MC3 when the time limit is higher. This suggests MC3 is better suited for quickly solving easier instances while DS is better for harder instances. To get a better understanding of this behavior, the right pane of Fig. 4 depicts the number of the high-level iterations of AA-CCBS with MC3, DS, DS+MC3 for each MAPF instance that was solved by all these methods. The instances are sorted by the number of iterations of MC3 (so its line monotonically increases) and the $x$-axis corresponds to distinct instances. As can be seen, in nearly half of the instances MC3 requires fewer high-level iterations than DS and DS+MC3. These are the cases that in general required fewer high-level iterations. Overall, our results confirm that both MC and DS are valuable techniques to enhance AA-CCBS performance. MC3 is well-suited for the easy instances while DS is better for the harder ones.
A comparison of AA-CCBS with its non any-angle variant is presented in Appendix.
@@ VII Bounded-Suboptimal AA-CCBS
Given a real number $w>1$, a bounded suboptimal (BS) solution is a solution whose cost does not exceed the cost of an optimal solution by more than a factor of $w$. Finding a BS solution is often easier than finding an optimal one. Creating a BS variant of AA-CCBS with any of our enhancements can be done by replacing the best-first search in the high-level of AA-CCBS with a Focal Search [33]. [^2] Creating a BS low-level search is not trivial in is beyond the scope of the paper. In more detail, in each high-level iteration we create a list of nodes called the FOCAL. FOCAL contains all the generated nodes whose cost does not exceed $w\cdot f_{min}$, where $f_{min}$ is the cost of the node regular AA-CCBS would have expanded. Our BS version of AA-CCBS chooses to expand the node in FOCAL with the smallest number of conflicts, breaking ties in favor of nodes with more constraints. This forces the search to resolve conflicts more greedily, while maintaining the required suboptimality bound.
We implemented this BS version of AA-CCBS with the three best-performing enhancements, DS, MC3, DS+MC3, and evaluated them on the same problem instances as before. Fig. 5 shows experimental results over the same problem instances as above across all maps using our BS version of AA-CCBS for suboptimality bounds $w=1.01$, $1.1$, and $1.25$ with the three best-performing enhancements, DS, MC3, DS+MC3. Fig. 5(top) depicts a box-and-whiskers plot of the solution cost overhead w.r.t. the SOC of the optimal solution. Fig. 5(bottom) shows the number of solved instances within our 300 seconds time limit (coverage).
In terms of coverage, MC3 notably increases the performance of AA-CCBS when searching for bounded suboptimal solutions (i.e., when $w>1$), while DS variants are not even able to outperform vanilla AA-CCBS. The poor results of DS may be due to the positive constraints it imposes, which often do not lead to immediately eliminating the conflict that causes these constraints. When searching for optimal solutions this is compensated by reducing the branching factor. However, the search for suboptimal solutions is more greedy by design and the branching factor becomes much less important. Thus, DS loses much of its power, while retaining the drawback of not immediately eliminating the conflict. At the same time, MC3 imposes constraints on a larger number of actions, and as the search greedily goes deeper it actually eliminates the conflict.
Moreover, due to the requirement to satisfy all the imposed positive constraints, the actual costs of the solutions found by DS versions are significantly higher than the ones of MC3. This can be seen in Fig. 5(top). The actual overhead in solution costs of MC3 version is not higher than 2% on average even when the suboptimality bound is $w=1.25$. The obtained results provide further support that MC3 is a useful technique: not only does it speed the search for optimal solutions on easy instances, but it also significantly increases the performance of bounded-suboptimal AA-CCBS.
@@ VIII CONCLUSION AND FUTURE WORK
ACKNOWLEDGMENT
This research is partially funded by ISF grant #1238/23 to Roni Stern and by the Ministry of Science and Higher Education of the Russian Federation (Project No. 075-15-2024-544).
APPENDIX
VIII-A Additional Experiments
To better understand what is the effect of allowing any-angle moves for the agents in MAPF we have compared AA-CCBS with its non-any-angle counterpart, i.e. with the original CCBS solver that allowed moves into 4 cardinal directions on the grid.
For this experiment we have used the most advanced version of CCBS described in [31]. It leverages such enhancements as prioritizing conflicts, disjoint-splitting and high-level heuristics. We denote this version as CCBS (cardinal-only). We compare it with the two variants of the proposed any-angle MAPF solver, i.e. AA-CCBS+DS+MC3 and AA-CCBS+MC3 with the suboptimality factor $w=1.25$. These solvers are denoted as AA-CCBS (optimal) and AA-CCBS (suboptimal).
The evaluation protocol and setup was the same as in the experiments reported in the main body of the paper (we used the same maps, problem instances, time cap etc.). The results are reported in Table I and Fig. 6.
CCBS AA-CCBS AA-CCBS
(cardinal-only) (optimal) (suboptimal)
empty-16-16 127.4% 100.0% 101.4%
random-32-32-20 117.6% 100.0% 101.1%
maze-32-32-4 115.6% 100.0% 100.4%
den312d 117.9% 100.0% 100.2%
warehouse-10-20-10-2-2 111.8% 100.0% 100.2%
TABLE I: The solution cost of CCBS (cardinal-only) compared to AA-CCBS (optimal) and AA-CCBS (suboptimal). The cost of AA-CCBS (optimal) is taken for 100%.
Table I shows the average solution cost relative to AA-CCBS (optimal). Recall that CCBS is an optimal solver so its solutions (under the given movement assumption) can not be improved. Still, as one can see from the table, the gap in solution cost between AA-CCBS (optimal) and CCBS (cardinal-only) can reach $27.4\%$. The lowest difference in solution cost is observed on the warehouse map. This can be explained by the structure of this map that consists of numerous narrow corridors that limit the ability to perform any-angle moves. When the map contains empty areas the gap is getting bigger.
Figure 6 demonstrates the success rate of the evaluated solvers across all maps. X-axis of each plot shows the number of agents and Y-axis shows the percentage of the problem instances (with this number of agents) successfully solved by a solver within a given time limit (of 300 seconds).
As CCBS (cardinal only) does not consider any-angle moves its search space is much more compact. Thus it manages to consistently outperform both optimal and suboptimal variants of AA-CCBS in terms of success rate. It is worth noting that AA-CCBS (suboprimal) performs notably better than AA-CCBS (optimal) and, generally, can be considered to provide a good trade off between the solution cost and planning time.
Figure 6: Success rates per number of agents for AA-CCBS and CCBS on different maps.